OpenStudy (anonymous):

Using complete sentences, explain the two ways to simplify sqrt(147) and provide the simplified form. You can show a square root by writing "sqrt()". For example, can be written as "sqrt(2)".

OpenStudy (raden):

setting 147 = 3 * 49 so, sqrt(147) = sqrt(3 * 49) = sqrt(3) * sqrt(49) = 7 sqrt)3) does that make sense ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but what about another way

OpenStudy (raden):

nothing :)

OpenStudy (anonymous):

what do you mean is that the only way

OpenStudy (raden):

ah, maybe like this sqrt(147) = sqrt(3 * 49) = sqrt(3) sqrt(49) = sqrt(3) (49)^1/2 = sqrt(3) (7^2)^1/2 = sqrt(3) 7^(2 * 1/2) = 7 sqrt(3) a bit different, heh :)

OpenStudy (anonymous):

yes

OpenStudy (raden):

but the result is same :)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

thanks

OpenStudy (raden):

you're welcome